# Precise Measurements of the LMC Bar’s Geometry With Gaia DR3 and a Novel Solution to Crowding Induced Incompleteness in Star Counting

### v1

@@ Abstract
We present new measurements of the two-dimensional (2-D) geometry of the LMC’s stellar bar with precise astrometric observations of red clump stars in Gaia DR3. We develop a novel solution to tackle crowding induced incompleteness in Gaia datasets with the Gaia BP-RP color excess. Utilizing the color excess information, we derive a 2-D completeness map of the LMC’s disk. We find that incompleteness biases the bar measurements and induces large uncertainties. With the completeness-corrected 2-D red clump map, we precisely measure the LMC bar’s properties for the first time using Fourier decomposition. The bar radius is $R_{bar}=2.13^{+0.03}_{-0.04}$ kpc, and its position angle is $121.26^{\circ}\pm 0.21^{\circ}$. The bar’s strength as quantified by the Fourier bi-symmetric amplitude is $S_{bar}=0.27$, indicating that the LMC has a significant bar perturbation. We find the bar has an axis ratio of $0.54\pm 0.03$, and is offset with respect to the center of the outer disk isophote at R $\approx$ 5 kpc by $0.76\pm 0.01$ kpc. These LMC bar properties agree with a hydrodynamic model where the SMC has undergone a recent direct collision with the LMC. We compare the LMC’s bar properties with other barred galaxies in the local universe, and discover that the LMC is similar to other barred galaxies in terms of bar-galaxy scaling relations. We discuss how our completeness correction framework can be applied to other systems in the Local Group.
facilities: Gaia, Astro Data Lab, UArizona HPCsoftware: This work made use of python, and its packages like numpy (van der Walt et al., 2011; Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007) and astropy (Astropy Collaboration et al., 2013, 2018, 2022). This work made use of the open source database management software topcat [^4] https://www.star.bris.ac.uk/mbt/topcat/(Taylor, 2005).
@@ 1 Introduction
The stellar bar of the LMC is one of the most peculiar features of this massive Milky Way satellite galaxy. Unlike bars in most galaxies of the local volume, the LMC’s bar is offset from the isophotal center of the disk (de Vaucouleurs & Freeman, 1972; van der Marel, 2001), is absent in the LMC’s interstellar medium (ISM) (Staveley-Smith et al., 2003), and is not evident in maps of young stellar populations (Harris & Zaritsky, 2009; Mazzi et al., 2021). The nature of the LMC’s bar is a topic of debate. Precise characterization of the geometry and strength of the LMC’s bar is needed to place the LMC in context with other barred galaxies and understand the role of bars in general in the evolution of low mass galaxies.
@@
As distance measurements have become more accurate, a different picture has emerged. Using distances to red clump stars, it has been found that the bar actually resides in the disk plane (Subramaniam & Subramanian, 2009; Haschke et al., 2012). Using state-of-the-art observations from the Survey of MAgellanic Stellar History (SMASH Nidever et al., 2017), Choi et al. (2018a) also found that the bar is consistent with being in the disk plane, but is likely tilted with respect to the disk by as much as $\sim 15^{\circ}$. The observation that the bar is tilted helps to understand how earlier studies could conclude that the LMC’s bar was not in the disk plane.
@@
In addition to the nature of the LMC’s stellar bar, the connection between the bar and surrounding ISM is not well understood. HI density and velocity maps of the LMC indicate no signature of a bar (Staveley-Smith et al., 2003), which is in contrast to what is observed for most massive barred galaxies where the bar is evident both in HI and other ISM tracers (e.g. Bosma, 1978; Fathi et al., 2005; López-Cobá et al., 2022). The absence of the LMC’s bar in its ISM gives credence to the idea that the LMC bar is not a standard dynamical bar, making the need to accurately characterize the LMC bar’s properties even more important.
Various attempts have been made to measure the 2-D geometry of the LMC’s stellar bar, which is quantified by the bar position angle, radius, ellipticity (axis ratio), and offset of the bar center with respect to the center of the stellar disk isophotes. Table 1 provides a survey of estimates of these properties of the LMC’s bar in literature. The table illustrates that there is little consistency among different works. For example, the radius of the bar differs between studies by as much as a factor of two. Such a high uncertainty is also an impediment towards understanding the origin of the peculiar properties of the bar. For example, an accuracy of at least 0.5 kpc is needed for parameters like the bar radius to accurately determine the torques on the bar from external perturbers like the SMC (H. Rathore et al. 2024(b), in prep).
@@ 0.87 Zhao & Evans (2000)
  • Incomplete stellar photometry due to crowding: the stellar field in the central region of the LMC is highly crowded, which significantly affects measurements of the bar structure that are based on resolved stars (e.g., Choi et al., 2018a). Crowding manifests as less complete stellar photometry in the central regions of the LMC, which leads to a systematic underestimate of the stellar density field in the bar region. Incomplete stellar photometry prevents an accurate measurement of the dynamical strength of the bar, which particularly hinders the comparison of the LMC to other barred galaxies. To the best of our knowledge, the LMC bar’s strength has not been measured before.
@@ 2 Gaia DR3 Data for the LMC and a Novel Solution to Incompleteness
Gaia has ushered a new era of galaxy dynamics by providing precise astrometry for stars in the Milky Way and nearby galaxies like the LMC (e.g. Gaia Collaboration et al. (2021) and references therein). These data bring novel opportunities for precise characterization of the structure and kinematics of the LMC, including its stellar bar (e.g. Zivick et al., 2019; Choi et al., 2022; Dhanush et al., 2024; Jiménez-Arranz et al., 2024).
We utilize red clump stars in Gaia DR3 (Gaia Collaboration et al., 2016; Lindegren et al., 2021; Gaia Collaboration et al., 2023) for precise geometric characterization of the LMC’s bar. The LMC’s bar predominantly consists of older stellar populations (Harris & Zaritsky, 2009; Mazzi et al., 2021), thus red clump stars are representative of the bar. Moreover, red clump stars occupy a narrow and a well defined region in the Hertzsprung-Russel diagram (Castellani et al., 2000), which makes them easy to select in Gaia Color-Magnitude diagrams (e.g. Choi et al., 2022). Using red clump stars will also allow us to place our results in context with several other works that have utilized this stellar population to study the LMC’s bar and disk (e.g. Subramaniam & Subramanian, 2009; Choi et al., 2018a, 2022). Red clump stars have a similar mass (Castellani et al., 2000), which makes comparisons with numerical simulations where star particles have equal masses easier. Further, red clump stars are standard candles for distance estimation (Girardi, 2016), offering potential for extending our framework to determine the 3-D geometry of the LMC’s bar in future analyses.
We follow the selection criteria of Choi et al. (2022) (hereafter C22) for selecting red clump stars in the LMC using the Gaia query interface  [^1] https://gea.esac.esa.int/archive/. The primary difference between our selection and the selection of C22 is that we are using Gaia DR3, whereas C22 used Gaia EDR3. Gaia DR3 offers a more extensive sample of stars with reduced errors compared to EDR3 (Jiménez-Arranz et al., 2023). We describe the exact Gaia query and selection procedure for the red clump stars in Appendix A. At this stage, we have a sample of 1,051,494 red clump stars. We refer to this sample as our parent sample. We utilize the neural network based catalog of Jiménez-Arranz et al. (2023) to evaluate the contamination of Milky Way foreground stars in our sample of LMC stars. We estimate this contamination to be less than 1% in the LMC bar region (R $<3^{\circ}$).
@@ 2.1 Coordinate Systems
For analyzing the geometry of the LMC’s bar and for comparing it with numerical simulations, we use the LMC cartesian coordinate system (x [kpc], y [kpc]). This coordinate system is obtained by defining an infinitely thin plane across the LMC’s disk that is inclined to the sky by the LMC’s observed inclination about its line of nodes and oriented at the observed position angle of the line of nodes.
We describe how to obtain these coordinates from RA - DEC coordinates in Appendix B.
@@ 2.2 Motivating an Incompleteness Correction Procedure Using the Gaia Color Excess
The flux measurement process for the G filter is different from the BP and RP filters. The flux in the G filter is obtained using profile fitting to an image with a resolution of $0.4$ arcsec, whereas the fluxes in BP and RP filters are obtained with an aperture of size $3.5\times 2.1$ arcsec2 (Evans et al., 2018). In crowded fields, fluxes within the BP and RP apertures can be contaminated by nearby sources. This contamination will in general be stochastic, and different for each star in the crowded field. Thus, the scatter in $C^{\ast}$ at a given location in the LMC’s disk probes the extent of crowding in that location.
@@ 3 Deriving a Completeness Map for the LMC
In the left panel of Figure 2, we show the spatial distribution of $\sigma_{C^{\ast}}$, in bins of $(\Delta\xi\times\Delta\eta)=(0.02^{\circ}\times 0.02^{\circ})$. We observe that $\sigma_{C^{\ast}}$ traces crowded regions in the LMC’s disk. However, $C^{\ast}$ may have an intrinsic scatter ($\sigma_{C^{\ast}}^{int}$) that is related to instrumental and astrophysical [^2] For example, for completely isolated sources, the combined BP+RP flux compared to the G flux may be slightly different from source to source because of differences in their spectral energy distributions (SEDs). effects, but not to stellar crowding itself. We measure $\sigma_{C^{\ast}}^{int}$ by analyzing the spatial distribution of $C^{\ast}$ in the outskirts of the LMC’s disk, where crowding does not significantly affect stellar photometry. In the middle panel of Figure 2, we show the radial profile of the spatial distribution of $\sigma_{C^{\ast}}$ and find that indeed the profile flattens out to a constant value at larger radii. To determine this constant value, we fit the radial profile with a tangent-hyperbolic model:
@@
We subtract the intrinsic scatter in $C^{\ast}$ from the overall scatter in quadrature to obtain a statistical quantity that probes crowding, which we call $\sigma_{C^{\ast}}^{crowd}$. We find that the dependence of $\sigma_{C^{\ast}}$ on other quantities like color and magnitude is less than its intrinsic scatter, and thus is not expected to affect the completeness calibration. Hence, we consider just the spatial dependence.
@@
Figure 2: Left Panel: The spatial distribution of the spread in the corrected Gaia color excess($\sigma_{C^{\ast}}$), which traces crowded regions in the LMC’s disk. Middle Panel: The radial profile of the spatial distribution of $\sigma_{C^{\ast}}$. The profile approaches a constant value in the outer disk, which indicates an intrinsic scatter to $C^{\ast}$ ($\sigma_{C^{\ast}}^{int}$) presumably due to instrumental effects independent of crowding. We find this constant value to be 0.04 (depicted by the black dashed line) by fitting the radial profile data with a tangent hyperbolic function. Right Panel: The spatial distribution of $\sigma_{C^{\ast}}^{crowd}$, which we obtain by subtracting the intrinsic scatter in $C^{\ast}$ from the total scatter in quadrature. $\sigma_{C^{\ast}}^{crowd}$ directly probes crowding in the LMC’s disk.
@@
$f=f_{0}\left[1-\tanh{\left(\frac{\sigma_{C^{\ast}}^{crow}-\sigma_{0}}{\sigma_{s}}\right)}\right]$ (5)
@@
We find that even after the completeness correction, there exist a few patches in the LMC’s disk where the star counts drop locally. We attribute these patches to regions where the dust extinction is high. Our framework will miss red clump stars that reside towards the high-end tail of the reddening distribution, even after the completeness correction, since they are omitted from our simple red clump selection box. We currently do not have a prescription to correct the counts for this effect. However, our completeness correction fully accounts for the selected red clump stars, so we do not expect the missed, highly-extincted red clump stars to affect our bar analysis.
@@
Figure 4: Completeness corrected LMC disk obtained by weighting the distribution in Figure 3 by the inverse of the completeness map. Central region becomes much more prominent after the completeness correction.
@@
Jiménez-Arranz et al. (2024) (hereafter JA24) use Gaia selection functions (Cantat-Gaudin et al., 2023; Castro-Ginard et al., 2023) to estimate an average completeness $>$ 50% in the LMC bar region for stars with $19.519<G<$ 19.5. Based on our color excess framework, we obtain a median completeness of 64% for the red clump stars ($19.318.6<G<$ 19.3) within $2.5^{\circ}$ of the LMC’s center, which is consistent with the prediction of JA24 for our red clump magnitude range. This higher completeness compared to JA24 is expected because our analysis uses brighter stars.
@@ 4.2 Centering the LMC’s Disk
In Table 2, we report our stellar dynamical center of mass estimate for the LMC for both incomplete and completeness corrected samples, along with other center of mass values reported in literature. In Figure 6, we show the location of the various center of masses. We find that incompleteness significantly affects the inference of the dynamical center of mass.
@@
Dynamical (Completeness Corrected) This Work (80.27, -69.65) 0
Dynamical (Incomplete) This work (78.89, -70.04) 0.62
Kinematic C22 (80.44, -69.27) 0.38

### v2

@@ Abstract
We present new measurements of the two-dimensional (2-D) geometry of the LMC’s stellar bar with precise astrometric observations of red clump stars in Gaia DR3. We develop a novel solution to tackle crowding induced incompleteness in Gaia datasets with the Gaia BP-RP color excess. Utilizing the color excess information, we derive a 2-D completeness map of the LMC’s disk. We find that incompleteness biases the bar measurements and induces large uncertainties. With the completeness-corrected 2-D red clump map, we precisely measure the LMC bar’s properties using Fourier decomposition. The bar radius (semi-major axis) is $R_{bar}=2.13^{+0.03}_{-0.04}$ kpc, and its position angle is $121.26^{\circ}\pm 0.21^{\circ}$. The bar’s strength as quantified by the Fourier bi-symmetric amplitude is $S_{bar}=0.27$, indicating that the LMC has a significant bar perturbation. We find the bar has an axis ratio of $0.54\pm 0.03$, and is offset with respect to the center of the outer disk isophote at R $\approx$ 5 kpc by $0.76\pm 0.01$ kpc. These LMC bar properties agree with a hydrodynamic model where the SMC has undergone a recent direct collision with the LMC. We compare the LMC’s bar properties with other barred galaxies in the local universe, and discover that the LMC is similar to other barred galaxies in terms of bar-galaxy scaling relations. We discuss how our completeness correction framework can be applied to other systems in the Local Group.
facilities: Gaia, Astro Data Lab, UArizona HPCsoftware: This work made use of python, and its packages like numpy (van der Walt et al., 2011; Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007) and astropy (Astropy Collaboration et al., 2013, 2018, 2022). This work made use of the open source database management software topcat (Taylor, 2005).
@@ 1 Introduction
The stellar bar of the LMC is one of the most peculiar features of this massive Milky Way satellite galaxy. Unlike bars in most gas rich galaxies of the local volume, the LMC’s bar is offset from the isophotal center of the disk (de Vaucouleurs & Freeman, 1972; van der Marel, 2001) and is absent in the LMC’s interstellar medium (ISM) (Staveley-Smith et al., 2003). Moreover, the bar is significantly more dominant in the older stellar populations (age $>$ 1 Gyr) as compared to the younger stellar populations (age $<$ 100 Myr) (Harris & Zaritsky, 2009; El Youssoufi et al., 2019; Mazzi et al., 2021; Gaia Collaboration et al., 2021). The nature of the LMC’s bar is a topic of debate. Precise characterization of the geometry and strength of the LMC’s bar is needed to place the LMC in context with other barred galaxies in the local universe and understand the role of bars in general in the evolution of low mass galaxies.
@@
As distance measurements have become more accurate, a different picture has emerged. Using distances to red clump stars, it has been found that the bar actually resides in the disk plane (Subramaniam & Subramanian, 2009). Using state-of-the-art observations from the Survey of MAgellanic Stellar History (SMASH Nidever et al., 2017), Choi et al. (2018a) found that the bar is consistent with being in the disk plane, but is likely tilted with respect to the disk by as much as $\sim 15^{\circ}$. Haschke et al. (2012) also infer that the LMC’s bar resides in a different plane relative to the disk. The observation that the bar is tilted helps to understand how earlier studies could conclude that the LMC’s bar was not in the disk plane.
@@
In addition to the nature of the LMC’s stellar bar, the connection between the bar and surrounding ISM is not well understood. HI density and velocity maps of the LMC indicate no signature of a bar (Staveley-Smith et al., 2003), which is in contrast to what is observed for most massive gas-rich barred galaxies where the bar is evident both in HI and other ISM tracers (e.g. Bosma, 1978; Fathi et al., 2005; López-Cobá et al., 2022), even when the bar is offset from the disk center (Colin & Athanassoula, 1989). The absence of the LMC’s bar in its ISM gives credence to the idea that the LMC bar is not a standard dynamical bar, making the need to accurately characterize the LMC bar’s properties even more important.
Various attempts have been made to measure the 2-D geometry of the LMC’s stellar bar, which is quantified by the bar position angle, radius (semi-major axis), ellipticity (axis ratio), and offset of the bar center with respect to the center of the stellar disk isophotes. Table 1 provides a survey of estimates of these properties of the LMC’s bar in literature. The table illustrates that there is little consistency among different works. For example, the radius of the bar differs between studies by as much as a factor of two. Such a high uncertainty is also an impediment towards understanding the origin of the peculiar properties of the bar. For example, an accuracy of at least 0.5 kpc is needed for parameters like the bar radius to accurately determine the torques on the bar from external perturbers like the SMC (H. Rathore et al. 2024(b), in prep).
@@ 0.87 Zhao & Evans (2000)
  • Incomplete stellar photometry due to crowding: the stellar field in the central region of the LMC is highly crowded, which significantly affects measurements of the bar structure that are based on resolved stars (e.g., Choi et al., 2018a). Crowding manifests as less complete stellar photometry in the central regions of the LMC, which leads to a systematic underestimate of the stellar density field in the bar region. Incomplete stellar photometry prevents an accurate measurement of the dynamical strength of the bar, which particularly hinders the comparison of the LMC to other barred galaxies. To the best of our knowledge, the LMC bar’s strength has not been measured before with the LMC’s stellar density distribution.
@@ 2 Gaia DR3 Data for the LMC and a Novel Solution to Incompleteness
Gaia has ushered a new era of galaxy dynamics by providing precise astrometry for stars in the Milky Way and nearby galaxies like the LMC (e.g. Gaia Collaboration et al. (2021) and references therein). These data bring novel opportunities for precise characterization of the structure and kinematics of the LMC, including its stellar bar (e.g. Zivick et al., 2019; Choi et al., 2022; Dhanush et al., 2024; Jiménez-Arranz et al., 2024b).
We utilize red clump stars in Gaia DR3 (Gaia Collaboration et al., 2016; Lindegren et al., 2021; Gaia Collaboration et al., 2023) for precise geometric characterization of the LMC’s bar. The LMC’s bar predominantly consists of older stellar populations (Harris & Zaritsky, 2009; Mazzi et al., 2021; Gaia Collaboration et al., 2021), thus red clump stars are representative of the bar. Moreover, red clump stars occupy a narrow and a well defined region in the Hertzsprung-Russel diagram (Castellani et al., 2000), which makes them easy to select in Gaia Color-Magnitude diagrams (e.g. Choi et al., 2022). Using red clump stars will also allow us to place our results in context with several other works that have utilized this stellar population to study the LMC’s bar and disk (e.g. Subramaniam & Subramanian, 2009; Choi et al., 2018a, 2022). Red clump stars have a similar mass (Castellani et al., 2000), which makes comparisons with numerical simulations where star particles have equal masses easier. Further, red clump stars are standard candles for distance estimation (Girardi, 2016), offering potential for extending our framework to determine the 3-D geometry of the LMC’s bar in future analyses.
We follow the selection criteria of Choi et al. (2022) (hereafter C22) for selecting red clump stars in the LMC using the Gaia query interface  [^1] https://gea.esac.esa.int/archive/. We describe the exact Gaia query and selection procedure for the red clump stars in Appendix A. At this stage, we have a sample of 1,051,494 red clump stars. We refer to this sample as our parent sample. We utilize the neural network based catalog of Jiménez-Arranz et al. (2023) to evaluate the contamination of Milky Way foreground stars in our sample of LMC stars. We estimate this contamination to be less than 1% in the LMC bar region (R $<3^{\circ}$) for their NN complete sample.
@@ 2.1 Coordinate Systems
For analyzing the geometry of the LMC’s bar and for comparing it with numerical simulations, we use the LMC in-plane cartesian coordinate system (x [kpc], y [kpc]). This coordinate system is obtained by defining an infinitely thin plane across the LMC’s disk that is inclined to the sky by the LMC’s observed inclination about its line of nodes and oriented at the observed position angle of the line of nodes.
We describe how to obtain these coordinates from the RA - DEC coordinates in Appendix B.
@@ 2.2 Motivating an Incompleteness Correction Procedure Using the Gaia Color Excess
The flux measurement process for the G filter is different from the BP and RP filters. The flux in the G filter is obtained using profile fitting to an image with a resolution of $0.4$ arcsec, whereas the fluxes in BP and RP filters are obtained with an effective aperture of size $3.5\times 2.1$ arcsec2 (Evans et al., 2018). In crowded fields, fluxes within the BP and RP apertures can be contaminated by nearby sources. This contamination will in general be stochastic, and different for each star in the crowded field. Thus, the scatter in $C^{\ast}$ at a given location in the LMC’s disk probes the extent of crowding in that location.
@@ 3 Deriving a Completeness Map for the LMC
In the left panel of Figure 2, we show the spatial distribution of $\sigma_{C^{\ast}}$, in bins of $(\Delta\xi\times\Delta\eta)=(0.02^{\circ}\times 0.02^{\circ})$. We observe that $\sigma_{C^{\ast}}$ traces crowded regions in the LMC’s disk. However, $C^{\ast}$ may have an intrinsic scatter ($\sigma_{C^{\ast}}^{int}$) that is related to instrumental and astrophysical [^2] For example, for completely isolated sources, the combined BP+RP flux compared to the G flux may be slightly different from source to source because of differences in their spectral energy distributions (SEDs). effects, but not to stellar crowding itself. We measure $\sigma_{C^{\ast}}^{int}$ by analyzing the spatial distribution of $C^{\ast}$ in the outskirts of the LMC’s disk, where crowding does not significantly affect the stellar photometry. In the middle panel of Figure 2, we show the radial profile of the spatial distribution of $\sigma_{C^{\ast}}$ and find that indeed the profile flattens out to a constant value at larger radii. To determine this constant value, we fit the radial profile with a tangent-hyperbolic model:
@@
We subtract the intrinsic scatter in $C^{\ast}$ from the overall scatter in quadrature to obtain a statistical quantity that probes crowding, which we call $\sigma_{C^{\ast}}^{crowd}$. We find that the dependence of $\sigma_{C^{\ast}}$ on other quantities like color and magnitude is less than its intrinsic scatter, and thus is not expected to significantly affect the completeness calibration. Hence, we consider just the spatial dependence.
@@
Figure 2: Left Panel: The spatial distribution of the spread in the corrected Gaia color excess ($\sigma_{C^{\ast}}$), which traces crowded regions in the LMC’s disk. Middle Panel: The radial profile of the spatial distribution of $\sigma_{C^{\ast}}$. The profile approaches a constant value in the outer disk, which indicates an intrinsic scatter to $C^{\ast}$ ($\sigma_{C^{\ast}}^{int}$) presumably due to instrumental and astrophysical effects independent of crowding. We find this constant value to be 0.04 (depicted by the black dashed line) by fitting the radial profile data with a tangent hyperbolic function. Right Panel: The spatial distribution of $\sigma_{C^{\ast}}^{crowd}$, which we obtain by subtracting the intrinsic scatter in $C^{\ast}$ from the total scatter in quadrature. $\sigma_{C^{\ast}}^{crowd}$ directly probes crowding in the LMC’s disk.
@@
$f=f_{0}\left[1-\tanh{\left(\frac{\sigma_{C^{\ast}}^{crowd}-\sigma_{0}}{\sigma_{s}}\right)}\right]$ (5)
@@
We find that even after the completeness correction, there exist a few patches in the LMC’s disk where the star counts drop locally. We attribute these patches to regions where the dust extinction is high. Our framework will miss red clump stars that reside towards the high-end tail of the reddening distribution, even after the completeness correction, since they are omitted from our simple red clump selection box. We currently do not have a prescription to correct the counts for this effect. However, our completeness correction fully accounts for the selected red clump stars, so we do not expect the missed, highly-extincted red clump stars to significantly affect our bar analysis.
@@
Figure 4: Completeness corrected LMC disk obtained by weighting the incomplete disk in Figure 1 (right panel) by the inverse of the completeness map (Figure 3). Central region becomes much more prominent after the completeness correction.
@@
Jiménez-Arranz et al. (2024b) (hereafter JA24) use Gaia selection functions (Cantat-Gaudin et al., 2023; Castro-Ginard et al., 2023) to estimate an average completeness $>$ 50% in the LMC bar region for stars with $19.519<G<$ 19.5. Based on our color excess framework, we obtain a median completeness of 64% for the red clump stars ($19.318.6<G<$ 19.3) within $2.5^{\circ}$ of the LMC’s center, which is consistent with the prediction of JA24 for our red clump magnitude range. This higher completeness compared to JA24 is expected because our analysis uses brighter stars.
@@ 4.2 Centering the LMC’s Disk
In Table 2, we report our stellar dynamical center of mass estimate for the LMC for both incomplete and completeness corrected samples, along with other center of mass values reported in literature. In Figure 6, we show the location of the various center of masses. We find that incompleteness significantly affects the inference of the stellar dynamical center of mass.
@@
Stellar Dynamical (Completeness Corrected) This Work (80.27, -69.65) 0
Stellar Dynamical (Incomplete) This work (78.89, -70.04) 0.62
Stellar Kinematic C22 (80.44, -69.27) 0.38